OpenStudy (anonymous):

Which is equal to (3 x^2 yz)(5x y^3 z^2)? (Points : 1) 8x3y4z3 15x2y3z2 15x3y4z3 45x3y4z3

OpenStudy (anonymous):

Which is equal to (3x2yz)(5xy3z2)? 8x^3 y^4 z^3 15x^2 y^3 z^2 15 x^3 y^4 z^3 45 x^3 y^4 z^3 I GOT C

OpenStudy (anonymous):

\[(3x^{2}yz)(5xy^{3}z^{2}) = 15x^{3}y^{4}z^{3}\]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

"c" because when multiplying variables you add the exponents

OpenStudy (anonymous):

Bananas multiply bananas, apples multiply apples and pears multiply pears. \[3*5 = 15\] \[x^{2}*x = x^{3}\] \[y*y^{3} = y^{4}\] \[z*z^{2} = z^{3}\] Combining everything from the right, you get the answer. \[15x^{3}y^{4}z^{3}\]

OpenStudy (anonymous):

yes what about 5x^2 / 5 x ^3 would the answer be 1x^-1?

OpenStudy (anonymous):

or just x^-1?

OpenStudy (anonymous):

\[\frac{ 5x^{2} }{ 5x^{3} } = \frac{ 1 }{ x }\] Again, apples divide apples and pears divide pears. \[\frac{ 5 }{ 5 }=1\] \[\frac{ x^{2} }{ x^{3} }=\frac{ x*x }{ x*x*x }=\frac{ 1 }{ x }\] Multiplying the apples times the pears, you get the answer, 1/x.

OpenStudy (anonymous):

1x^-1 is the same as x^-1, as multiplying anything by one does absolutely nothing. (6 times 1 is 6, 8 times 1 is 8, etc.)

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

so 6^-1 would be -6?

OpenStudy (anonymous):

No, for any number you use, elevating it to a negative number, makes it go from the numerator to the denominator, or backwards if it's already in the denominator. \[x^{-1} = \frac{ 1 }{ x }\] So 6^-1 would be 1/6.

OpenStudy (anonymous):

oh ok :)